# Decoding for Punctured Convolutional and Turbo Codes: A Deep Learning Solution for Protocols Compliance

### v1

@@ 5.1 Computational Complexity Analysis
Although the CNE decoder results in a higher computational cost compared to traditional Viterbi and BCJR algorithms, this overhead is offset by the reduced number of required decoding iterations and the substantial performance gains demonstrated in both AWGN and Rayleigh fading channels. Furthermore, the parallelizable nature of matrix operations in the LSTM and fully connected layers allows for efficient deployment on modern hardware accelerators, such as GPUs, thereby enhancing the overall computational efficiency.
5.2 Decoding Latency Analysis
Decoding latency is a key performance metric in practical communication systems, as it directly impacts throughput and real-time system performance. In this subsection, we analyze the decoding latency of the proposed LSTM-based CNE decoder and compare it with traditional Viterbi and BCJR algorithms, focusing on the computational delay of each module to provide a comprehensive understanding of their respective latency characteristics.
The decoding latency of the proposed CNE decoder stems from four primary components: the input projection and gating mechanism, the batch normalization, the LSTM network, and the output projection. The input features $\boldsymbol{x}\in\mathbb{R}^{K\times D_{\text{in}}}$ are first projected through a fully connected layer parameterized by $\boldsymbol{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$. Concurrently, a gating mechanism modulates these features based on the puncturing input $\boldsymbol{p}\in\mathbb{R}^{K\times D_{\text{in}}}$, using a gating weight matrix $\boldsymbol{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$. Both operations primarily involve matrix-vector multiplications, and assuming parallel processing across all $K$ time steps, the total latency of this stage can be expressed as:
$T_{\text{proj}}=t_{\text{mat}}(D_{\text{embed}},D_{\text{in}})$ (28)
where $t_{\text{mat}}(m,n)$ represents the processing delay for a matrix-vector multiplication of size $m\times n$.
The modulated features $\in\mathbb{R}^{K\times D_{\text{embed}}}$ then undergo batch normalization, which normalizes the features independently across the embedding dimension $D_{\text{embed}}$. Since batch normalization operates in parallel across all time steps and dimensions, the latency of this step is:
$T_{\text{BN}}=t_{\text{bn}}(D_{\text{embed}})$ (29)
where $t_{\text{bn}}(d)$ denotes the processing time required to normalize a vector of length $d$. Given its simplicity and element-wise operation, batch normalization contributes minimally to the overall latency.
The batch-normalized features are then passed through the LSTM network, which constitutes the core computational module of the CNE decoder. Due to the sequential nature of LSTMs, the latency of this stage depends on the sequence length $K$, the input embedding size $D_{\text{embed}}$, and the hidden state size $D_{\text{hidden}}$. For each time step, the LSTM processes input features, updates its hidden state, and computes gating activations. Assuming $t_{\text{lstm}}(D_{\text{hidden}},D_{\text{embed}})$ is the processing delay for one LSTM cell, the total latency for the LSTM is:
This sequential operation across $K$ time steps makes the LSTM the dominant factor in the total decoding latency.
Finally, the LSTM outputs $\boldsymbol{h}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$ are projected to the output space via a fully connected layer parameterized by $\boldsymbol{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$. As this operation is parallelizable across all $K$ time steps, its latency can be expressed as:
$T_{\text{out}}=t_{\text{mat}}(1,2D_{\text{hidden}})$ (31)
By combining the latencies of all these components, the total decoding latency of the CNE decoder can be expressed as:
$T_{\text{CNE}}=T_{\text{proj}}+T_{\text{BN}}+T_{\text{LSTM}}+T_{\text{out}}$ (32)
For comparison, the decoding latency of the Viterbi algorithm is determined by its forward recursion through the trellis structure. For a convolutional code with $2^{L-1}$ states, each time step requires updating path metrics for all states. Assuming $t_{\text{state}}(2^{L-1})$ represents the processing delay per state, the total latency of the Viterbi decoder is:
$T_{\text{Viterbi}}=K\cdot t_{\text{state}}(2^{L-1})$ (33)
As the Viterbi algorithm operates sequentially across $K$ time steps, its latency scales linearly with the sequence length. Although efficient, its reliance on a fixed trellis structure limits its flexibility in adapting to varying channel conditions.
The BCJR algorithm, which forms the basis of Turbo decoding, incurs significantly higher latency due to its bidirectional nature and iterative processing. Each decoding iteration involves both forward and backward recursions over the trellis, with $2^{L-1}$ states per recursion. Assuming $t_{\text{state}}(2^{L-1})$ is the processing delay for each recursion, the latency for a single iteration is:
$T_{\text{recursion}}=2\cdot K\cdot t_{\text{state}}(2^{L-1})$ (34)
and considering $N_{\text{iter}}$ iterations for the Turbo decoder, the total latency of the BCJR-based Turbo decoder is:
$T_{\text{BCJR}}=N_{\text{iter}}\cdot T_{\text{recursion}}=2\cdot N_{\text{iter}}\cdot K\cdot t_{\text{state}}(2^{L-1})$ (35)
[…]
In summary, the proposed LSTM-based CNE decoder strikes an effective balance between decoding performance and latency. While the sequential nature of the LSTM introduces higher latency compared to the Viterbi and BCJR algorithms, it outperforms under dynamic channel conditions and requires significantly fewer iterations than the BCJR-based Turbo decoder (e.g., 3 iterations versus 6 iterations in our experiments). Moreover, the CNE decoder benefits from parallelizable components such as input projection, batch normalization, and output projection, enabling efficient implementation on modern hardware accelerators. These advantages make the proposed CNE decoder highly suitable for real-time communication systems requiring both low latency and reliable performance.
@@ 6 Conclusions
This paper introduces a unified LSTM-based decoding architecture that enhances the performance of punctured convolutional and Turbo codes in practical communication scenarios. By leveraging deep learning techniques, the proposed method offers a flexible and code-agnostic solution that ensures robust decoding across a wide range of code rates and channel conditions. The results obtained from extensive simulations validate the efficacy of the approach, demonstrating notable improvements in decoding accuracy compared to traditional algorithms. Future work could focus on further optimizing the architecture to reduce decoding complexity and latency, as well as extending its application to more complex coding schemes and diverse communication environments, paving the way for more efficient and reliable decoding in next-generation AI-powered wireless systems.

### v3

@@ 5.5 Mitigating Complexity and Latency in Practice
  2. 2.
    Hardware Acceleration: Beyond GPUs, tailored FPGA-based accelerators offer a promising solution. Pipelined Vector-Scalar Multiplication engines minimize latency through streamlined computation [47]. On-chip static random access memory (SRAM) and embedded dynamic random-access memory (eDRAM) optimize data access, reducing delays critical for real-time systems [48, 49]. Furthermore, partial reconfiguration adapts hardware dynamically, balancing performance and power consumption for constrained environments [50].
  3. 3.
    Scheduling Efficiency: Advanced scheduling strategies enhance latency performance. Unfolded and intergate scheduling parallelize LSTM computations across processing elements, significantly reducing latency for time-sensitive tasks [51, 52]. Loop unrolling and pipelining in FPGA designs further eliminate dependencies, ensuring deterministic, low-latency performance [53].
This layered approach—optimizing the model, leveraging specialized hardware, and refining scheduling—effectively mitigates complexity and latency challenges. These strategies ensure the proposed CNE decoder achieves robust, real-time performance, even in resource-constrained settings, making it highly suitable for practical communication systems. Future work could explore specific case studies or quantitative evaluations of these techniques to further validate their effectiveness in diverse scenarios.
@@ 6 Conclusions
This paper introduces a unified LSTM-based decoding architecture that enhances the performance of punctured convolutional and Turbo codes in practical communication scenarios. By leveraging deep learning techniques, the proposed approach offers a flexible and code-agnostic solution that ensures robust decoding across a wide range of code rates and channel conditions. The results obtained from extensive simulations validate the efficacy of the approach, demonstrating notable improvements in decoding accuracy compared to traditional algorithms. Future work could focus on further optimizing the architecture to reduce decoding complexity and latency, as well as extending its application to more complex coding schemes and diverse communication environments, paving the way for more efficient and reliable decoding in next-generation AI-powered wireless systems.